OpenStudy (sissyedgar):

The total rainfall in Forks, Washington, this October was 8 3/8 in. In September, 60% less rain fell than in October. How much rain fell in September in Forks, Washington? Express your answer as a mixed number in simplest form. ( ) ( )-------in. ( )

OpenStudy (sissyedgar):

@Daniellelovee

OpenStudy (daniellelovee):

you have to do 83/8 - 60%

OpenStudy (sissyedgar):

33/52?

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 60}}\%\ of\ 8\frac{3}{8} \implies \cfrac{{\color{brown}{ 60}}}{100}\cdot 8\frac{3}{8}\implies {\color{brown}{ 60}}\div 100\cdot 8\frac{3}{8}\)

OpenStudy (daniellelovee):

oh I thought it was 

OpenStudy (sissyedgar):

:P

OpenStudy (sissyedgar):

im lost, the answer is....

OpenStudy (sissyedgar):

@jdoe0001

OpenStudy (sissyedgar):

@Daniellelovee

OpenStudy (sissyedgar):

im like so lost rn

OpenStudy (agl202):

5 1/40 hope it helped. Are u still doing the redo?

OpenStudy (sissyedgar):

ty

OpenStudy (sissyedgar):

redo?

OpenStudy (agl202):

Ya

OpenStudy (sissyedgar):

Arlo works at a paint store. A formula for a certain color calls for 3/5 L of green paint. Arlo wants to make 1/3 less paint than the formula creates. How much green paint should Arlo use? Express your answer as a fraction in simplest form. ( ) ------ L ( )

OpenStudy (sissyedgar):

last 1 ^.^

OpenStudy (agl202):

6/15

OpenStudy (agl202):

What did u get this time @sissyedgar ?

OpenStudy (sissyedgar):

lemme see

OpenStudy (sissyedgar):

uhhhhhhhhg!!!!!!!!!!!!!!!!

OpenStudy (sissyedgar):

I GOT 10%!!!

OpenStudy (agl202):

What ant the questions I answered?

OpenStudy (sissyedgar):

?

OpenStudy (agl202):

We're they correct?

OpenStudy (sissyedgar):

uh, i only got 1 right, at least one of the ones u gave me is wrong ill b bk later, i gotta go